# Electromagnetic response and emergent topological orders in transition metal dichalcogenide MoTe2 bilayers

### v1

@@ Appendix B Chern-Simons Flux attachment
We now relate the moiré filling $\nu$ to the filling of bands formed due to the interplay of the moiré potential and the CS field. For that, recall that a uniform field gives rise to Landau levels supporting $N_{LL}=B\,A_{system}/\phi_{0}$ states per LL, which can be re-expressed as $N_{LL}=B\,A_{system}/\phi_{0}=\frac{B\,A_{uc}}{\phi_{0}}\,N_{uc}=\frac{\phi}{\phi_{0}}\,N_{uc}=\frac{p}{q}N_{uc}$. Furthermore, we form the magnetic unit cell by enlarging the moiré unit cell by a factor of $q$, so that the number of states in the magnetic unit cell is $(q\,A_{uc})\,\rho_{LL}=(q\,A_{uc})\,\frac{1}{2\pi\ell^{2}_{B}}=(q\,A_{uc})\frac{1}{2\pi\frac{\hbar}{eB}}=(q\,A_{uc})\,\frac{B}{\phi_{0}}=q\,\frac{\phi}{\phi_{0}}=p$ That is, each LL gives rise to $p$ “subbands”, each of which supports a number of state
@@ Appendix C Quantum Phase Transitions
Due to the fractal nature of the moiré Hofstadter spectrum, we also observed quantum phase transitions at hole filling fractions $\nu_{h}=1/5$ and $2/9$, i.e., flux $\phi/\phi_{0}=2/5$ and $4/9$. It is worth pointing out that the filling fraction follows a $\frac{p}{4p+1}$ sequence.
@@ Appendix E Magnetic Translation Symmetry
As is well known, the presence of the magnetic field causes the Hamiltonian not to be invariant under regular translations. Let $W({\bm{a}})\equiv e^{i{\bm{p}}\cdot a/\hbar}$ be the generator of translations by a vector ${\bm{a}}$, such that $W({\bm{a}})\,{\bm{r}}\,W^{-1}({\bm{a}})={\bm{r}}+{\bm{a}}$ Then, $W({\bm{a}})\,{\bm{\Pi}}\,W^{-1}({\bm{a}})=\frac{1}{\hbar}({\bm{p}}+e{\bm{A}}({\bm{r}}+{\bm{a}}))=\frac{1}{\hbar}({\bm{p}}+e{\bm{A}}({\bm{r}})+e\,{\bm{\nabla}}f({\bm{r}}))$, where we have used the gauge equivalence ${\bm{A}}({\bm{r}}+{\bm{a}})={\bm{A}}({\bm{r}})+{\bm{\nabla}}f({\bm{r}})$ that is valid when the magnetic field is constant. The gradient of $f$ term can be “removed”, or gauged away, by a (unitary) gauge transformation
$e^{-i\,\frac{e}{\hbar}f({\bm{r}})}\,H[{\bm{A}}({\bm{r}})+{\bm{\nabla}}f]\,e^{i\,\frac{e}{\hbar}f({\bm{r}})}=.H[{\bm{A}}({\bm{r}})]\,.$ (45)
@@
$e^{-i\,\frac{e}{\hbar}f({\bm{r}})}\,W({\bm{a}})\,H[{\bm{A}}({\bm{r}})]\,W^{-1}({\bm{a}})\,,\,e^{i\,\frac{e}{\hbar}f({\bm{r}})}=.H[{\bm{A}}({\bm{r}})]\,.$ (46)

### v2

@@ Appendix B Chern-Simons Flux attachment
We now relate the moiré filling $\nu$ to the filling of bands formed due to the interplay of the moiré potential and the CS field. For that, recall that a uniform field gives rise to Landau levels supporting $N_{LL}=B\,A_{system}/\phi_{0}$ states per LL, which can be re-expressed as $N_{LL}=B\,A_{system}/\phi_{0}=\frac{B\,A_{uc}}{\phi_{0}}\,N_{uc}=\frac{\phi}{\phi_{0}}\,N_{uc}=\frac{p}{q}N_{uc}$. Furthermore, we form the magnetic unit cell by enlarging the moiré unit cell by a factor of $q$, so that the number of states in the magnetic unit cell is $(q\,A_{uc})\,\rho_{LL}=(q\,A_{uc})\,\frac{1}{2\pi\ell^{2}_{B}}=(q\,A_{uc})\frac{1}{2\pi\frac{\hbar}{eB}}=(q\,A_{uc})\,\frac{B}{\phi_{0}}=q\,\frac{\phi}{\phi_{0}}=p$ . That is, each LL gives rise to $p$ “subbands”, each of which supports a number of state
@@ Appendix C Quantum Phase Transitions
Due to the fractal nature of the moiré Hofstadter spectrum, we also observed quantum phase transitions at hole filling fractions $\nu_{h}=1/5$ and $2/9$, i.e., flux $\phi/\phi_{0}=2/5$ and $4/9$, as shown in Fig. 9 and Fig. 10. It is worth pointing out that the filling fraction follows a $\frac{p}{4p+1}$ sequence.
@@ Appendix E Magnetic Translation Symmetry
As is well known, the presence of the magnetic field causes the Hamiltonian not to be invariant under regular translations. Let $W({\bm{a}})\equiv e^{i{\bm{p}}\cdot a/\hbar}$ be the generator of translations by a vector $\bm{a}$}, such that $W({\bm{a}})\,{\bm{r}}\,W^{-1}({\bm{a}})={\bm{r}}+{\bm{a}}$ . Then, $W({\bm{a}})\,{\bm{\Pi}}\,W^{-1}({\bm{a}})=\frac{1}{\hbar}({\bm{p}}+e{\bm{A}}({\bm{r}}+{\bm{a}}))=\frac{1}{\hbar}({\bm{p}}+e{\bm{A}}({\bm{r}})+e\,{\bm{\nabla}}f({\bm{r}}))$, where we have used the gauge equivalence ${\bm{A}}({\bm{r}}+{\bm{a}})={\bm{A}}({\bm{r}})+{\bm{\nabla}}f({\bm{r}})$ that is valid when the magnetic field is constant. The gradient of $f$ term can be “removed”, or gauged away, by a (unitary) gauge transformation
$e^{-i\,\frac{e}{\hbar}f({\bm{r}})}\,H[{\bm{A}}({\bm{r}})+{\bm{\nabla}}f]\,e^{i\,\frac{e}{\hbar}f({\bm{r}})}=H[{\bm{A}}({\bm{r}})]\,.$ (45)
@@
$e^{-i\,\frac{e}{\hbar}f({\bm{r}})}\,W({\bm{a}})\,H[{\bm{A}}({\bm{r}})]\,W^{-1}({\bm{a}})\,,\,e^{i\,\frac{e}{\hbar}f({\bm{r}})}=H[{\bm{A}}({\bm{r}})]\,.$ (46)